OpenStudy (anonymous):

Charlotte has two job offers as a car sales person. The first job will pay her $600 a month plus 2% on the price of her car sales. The second job will pay $1000 a month plus 1% on the price of her sales. How much must she sell each month to make the same income from the two jobs? A. $20,000 B. $30,000 C. $40,000 D. $50,000

OpenStudy (anonymous):

I think it's C...but correct me if I'm wrong.

OpenStudy (mertsj):

600+.02x=1000+.01x

OpenStudy (mertsj):

Good for you. You got it right.

OpenStudy (ash2326):

Her income from first job= 600+0.02x x=price of her car sales Her income from second job=1000+0.01x We want to find x for which income from both the job is same 600+0.02x=1000+0.01x 0.01x=400 x=$40000. Kija you're correct. Great work:)

OpenStudy (anonymous):

YAY!!!! thnks u guys for checking for me.

OpenStudy (mertsj):

yw